# Comparing Two Proxy Methods for Causal Identification

### v2

@@ Abstract
Identifying causal effects in the presence of unmeasured variables is a fundamental challenge in causal inference, for which proxy variable methods have emerged as a powerful solution. We contrast two major approaches in this framework: (1) bridge equation methods, which leverage solutions to integral equations to recover causal targets, and (2) array decomposition methods, which recover latent factors composing counterfactual quantities by exploiting unique determination of eigenspaces. We compare the model restrictions underlying these two approaches and provide insight into implications of the underlying assumptions, clarifying the scope of applicability for each method.
@@ 1 Introduction
Identifying causal effects in the presence of unmeasured variables is a fundamental challenge in causal inference. Interest in proxy variable methods that leverage observed consequences of latent variables (i.e., proxies), has grown rapidly in both applied and methodological domains. Within this landscape, two foundational lines of work underpin nonparametric proxy variable methods for causal effect identification: (1) the bridge equation approach from Miao, Geng and Tchetgen Tchetgen (2018); and (2) the array decomposition approach, introduced to the causal literature by Kuroki and Pearl (2014) and Allman et al. (2015), building on ideas for full law identification in discrete latent variable models which trace back to Kruskal (1977). Both approaches admit generalizations, to varying degrees, to continuous settings.
@@
The bridge equation approach was developed for settings in which proxies are available for an unmeasured confounder, allowing the counterfactual distribution to be recovered directly through re-expression of the adjustment formula, without identification of the full law (Miao, Geng and Tchetgen Tchetgen, 2018). The method leverages solutions to bridge equations—a generalization of solutions to linear systems—to recover the causal target. This method was later adapted to settings where proxies are available for an unmeasured mediator in the absence of hidden confounder proxies, recovering causal effects through re-expression of the front-door adjustment formula (Ghassami et al., 2024).
@@
The array decomposition approach imposes similar invertibility conditions, but rather than solving directly for a causal target, it recovers the full law of latent and observed variables, from which causal effects follow. Originating from the work of Kruskal (1977), this framework allows causal effect identification in models with arbitrary hidden variables, such as in settings with unmeasured confounders (Kuroki and Pearl, 2014) or unmeasured exposures (Zhou and Tchetgen, 2024). Though initially formulated for effect identification in fully discrete models (Kuroki and Pearl, 2014; Allman et al., 2015), its identifying assumptions extend to continuous settings via the results of Hu and Schennach (2008), as shown by Deaner (2023).
Despite their shared ability to identify causal effects in the presence of unmeasured variables, these two approaches exhibit non-nested model restrictions. Prior work from Deaner (2023) has shown that the conditional independence assumptions (i.e., Markov restrictions) underlying the two approaches for hidden confounding are distinct and the models non-nested.
Until now, no work has characterized the relationship between the downstream assumptions required by each approach under shared conditional independences. This leaves unclear when a methodologist or practitioner should prefer one framework over the other. Hence, our work fills an important gap by providing a transparent comparison of the structural assumptions required by each method under shared Markov restrictions, clarifying how the two frameworks differ and the settings in which each is most appropriate.
@@ 1.1 Causal Graphical Models
Bayesian networks, represented as directed acyclic graphs (DAGs), offer a structured framework for statistical modeling, where nodes denote random variables and edges encode conditional independence assumptions (i.e., Markov restrictions) (Pearl, 1988). DAGs may be interpreted causally  (Pearl, 2009; Spirtes, Glymour and Scheines, 2001), encoding counterfactual relationships which enable formal reasoning about causal effects and their identifiability from observed data (Richardson and Robins, 2013). For example, Figure 1 depicts a simple confounding structure in which a common cause $C$ influences both treatment $A$ and outcome $Y$. This structure entails the counterfactual independence statement $Y(a)\perp\!\!\!\perp A\mid C$, known as conditional ignorability, where $Y(a)$ denotes a potential outcome random variable, interpreted to mean “$Y$ had $A$, possibly contrary to fact, been set to value aa”  (Richardson and Robins, 2013).
The methods we discuss rely on conditional independences and counterfactual assumptions reflected in causal graphical models referenced throughout the paper. In these figures, dashed edges ( ) indicate causal relationships that are optional and do not contradict identifying assumptions imposed.
@@ 1.2 Adjustment Formula
Consider the causal graphical model in Figure 1, where confounders $C$ between binary treatment $A$ and outcome $Y$ are fully observed. The counterfactual distribution $f_{Y(a)}(y)$ is identified by the adjustment formula (Equation 1) if the following hold for each $a\in\mathcal{A}$ and $c\in\mathcal{C}$:
@@
In general, Equation 1 holds if $C$ satisfies conditional ignorability assumption $Y(a)\perp\!\!\!\perp A\mid C$. This assumption is justified when, by the rules of d-separation, conditioning on $C$ blocks all backdoor paths (which start with an arrowhead into $A$) from $A$ to $Y$ (Pearl, 2009; Richardson and Robins, 2013).
@@
We provide an analogous review of the front-door adjustment formula in Appendix A.
@@ 2 Proximal Causal Learning via Bridge Equations
We now present the assumptions of the proximal causal learning approach via bridge equations, introduced by Miao, Geng and Tchetgen Tchetgen (2018) as a framework for addressing hidden confounding using confounder proxies, which enables recovery of the counterfactual distribution via the adjustment formula. Ghassami et al. (2024) develop analogous assumptions leveraging proxies of a hidden mediator in the absence of hidden confounder proxies, enabling identification causal effects via the frontdoor adjustment formula in Equation 7.
If confounders $C$ between treatment $A$ and outcome $Y$ are unobserved, the counterfactual distribution $f_{Y(a)}(y)$, and consequently average causal effects, are not identifiable without further assumptions.
Consider proxies $Z$ and $W$ for an unmeasured confounder (now denoted $U$), as illustrated in the causal graphical model in Figure 2. Note had $U$ been observed, the adjustment formula would yield the identifying expression:
$f_{Y(a)}(y)=\int f_{Y\mid U,A}(y\mid u,a)\,f_{U}(u)\,du.$ (3)
Miao, Geng and Tchetgen Tchetgen (2018) rewrite Equation 3 in terms of the observed law $f_{A,Y,W,Z}(a,y,w,z)$, leveraging the conditional independence restrictions listed in Assumption 4 and 5.
@@
In addition to Assumption 4 and 5, Miao, Geng and Tchetgen Tchetgen (2018) presume Assumptions 6 and 7, where $\mathbb{E}[g(u)\mid z,a]$ is a vector indexed by values $z\in\mathcal{Z}$. A detailed proof of identification is given in their paper. We will elaborate upon these additional two assumptions and their meanings in Sections 2.0.1 and 2.0.2 below.
@@
It is important to emphasize that while this approach identifies the interventional distribution $f_{Y(a)}(y)$, it does not, in general, identify factors of the full law of hidden and observed variables—for instance, the conditional distribution $f_{Y\mid U,A}(y\mid u,a)$ or the marginal distribution $f_{U}(u)$ of the unobserved variable. This stands in subtle contrast to the array decomposition approach discussed in Section 3.
2.0.1 Completeness Conditions
Assumption 6 generalizes the linear algebra notion of linear independence. In discrete models, Assumption 6 means that, for each $a\in\mathcal{A}$, the conditional laws $f_{Z\mid U,a}(z\mid u,a)$ span linearly independent directions across different values of $U$.
To illustrate this idea, suppose for the time being that the model depicted in Figure 2 is fully discrete. Define for each $a\in\mathcal{A}$,
$\mathbf{P}_{U\mid Z,a}=\big[\,f_{U\mid Z,A}(u\mid z_{1},a),\;\dots,\;f_{U\mid Z,A}(u\mid z_{|\mathcal{Z}|},a)\,\big],$
where columns are distribution vectors of the form
$f_{U\mid Z,a}(u\mid z,a)=\begin{bmatrix}\Pr(U=u_{1}\mid Z=z,A=a)\\ \vdots\\ \Pr(U=u_{|\mathcal{U}|}\mid Z=z,A=a)\end{bmatrix}$.
Assumption 6 is equivalent to right invertibility of matrix $\mathbf{P}_{U\mid Z,a}$. To see this, note the expansion of Assumption 6 below:
@@ 2.0.1 Completeness Conditions
Right invertibility of $\mathbf{P}_{U\mid Z,a}$ implies left invertibility of
$\mathbf{P}_{Z\mid U,a}=\big[\,f_{Z\mid U,A}(z\mid u_{1},a),\;\dots,\;f_{Z\mid U,A}(z\mid u_{|\mathcal{U}|},a)\,\big]$
since elementary row operations preserve matrix rank. Equivalently, for each $a\in\mathcal{A}$, the conditional laws $f_{Z\mid U,a}(z\mid u,a)$ span linearly independent directions across different values of $U$. Notably, $Z$ must have at least as many categories as $U$ in order for Assumption 6 to hold, as noted in Cui et al. (2023). Assumption 6 extends this idea to continuous settings.
Assumption 6 is often satisfied under widely used parametric specifications. For example, it holds when, for each treatment level $a$, the conditional distribution $f_{U\mid Z,a}(u\mid z,a)$ belongs to an exponential family indexed by $z$, provided that $z$ has support on an open interval of the real line (Brown, 1986; Newey and Powell, 2003). Since this setup may feel more natural when $Z$ acts as an instrumental variable rather than a proxy, it is worth noting that no identifying assumptions are violated under the modification $Z\to U$ in Figure 2. In addition, under standard conjugate–prior relationships for exponential families, $f_{U\mid Z,a}(u\mid z,a)$ may lie in the same exponential family as $f_{Z\mid U,a}(z\mid u,a)$.
Hu and Shiu (2022) provide several characterizations of nonparametric families of complete distributions, where $U$ is generated as an arbitrary function of $Z$ plus independent additive noise. In general, Assumption 6 is not testable without imposing further conditions (Canay, Santos and Shaikh, 2013).
2.0.2 Solution to a Bridge Equation
@@
The integral equation in Assumption 7, referred to as a bridge equation, enables recovery of the counterfactual distribution $f_{Y(a)}(y)$ alongside prior assumptions. Since the existence of a solution is usually assumed point blank, we unpack sufficient conditions for and implications of this assumption.
Define for each $a\in\mathcal{A}$,
$\mathbf{P}_{Y\mid Z,a}$ $=\big[\,f_{Y\mid Z,A}(y\mid z_{1},a),\;\dots,\;f_{Y\mid Z,A}(y\mid z_{|\mathcal{Z}|},a)\,\big],$
$\mathbf{P}_{W\mid Z,a}$ $=\big[\,f_{W\mid Z,A}(w\mid z_{1},a),\;\dots,\;f_{W\mid Z,A}(w\mid z_{|\mathcal{Z}|},a)\,\big].$
@@
In the discrete case, Assumption 7 is equivalent to the existence of a matrix $\mathbf{H}_{a}$ such that
$\mathbf{P}_{Y\mid Z,a}\;=\;\mathbf{H}_{a}\,\mathbf{P}_{W\mid Z,a}.$
A new completeness Assumption 7a shown below—i.e. that $\mathbf{P}_{W\mid Z,a}$ is left invertible in a discrete model—is sufficient for Assumption 7 in the discrete setting since
$\mathbf{P}_{Y\mid Z,a}=\mathbf{P}_{Y\mid Z,a}\,\mathbf{P}_{W\mid Z,a}^{\dagger}\,\mathbf{P}_{W\mid Z,a},$
where $\mathbf{P}_{W\mid Z,a}^{\dagger}$ denotes a left inverse of $\mathbf{P}_{W\mid Z,a}$. This requires $W$ to have at least as many categories as $Z$, and therefore at least as many as the latent variable $U$ under Assumption 6. From Assumption 6,
@@
Then taking the inner product with $f_{U}(u)$ on both sides gives $f_{Y(a)}(y)$.
In continuous models, Assumptions 7b7d in addition to Assumption 7a are sufficient to ensure that Assumption 7 holds (Miao, Geng and Tchetgen Tchetgen, 2018).
Assumption 7a.
For each $a\in\mathcal{A}$ and $g$ square-integrable,
$\mathbb{E}[g(z)\mid w,a]=0\,\text{ a.s.}\iff g(z)=0\,\text{ a.s.}$

### v3

@@ Abstract
Identifying causal effects in the presence of unmeasured variables is a fundamental challenge in causal inference, for which proxy variable methods have emerged as a powerful solution. We contrast two major approaches in this framework: (1) bridge equation methods, which leverage solutions to integral equations to recover causal targets, and (2) array decomposition methods, which recover latent factors used to identify counterfactual quantities via eigendecomposition tasks. We compare the model restrictions underlying these two approaches and provide insight into implications of the underlying assumptions, clarifying the scope of applicability for each method.
@@ 1 Introduction
Identifying causal effects in the presence of unmeasured variables is a fundamental challenge in causal inference. Interest in proxy variable methods, i.e. methods that leverage variables associated with latent variable(s), has grown rapidly in both applied and methodological domains. Within this landscape, two foundational lines of work underpin proxy-based approaches to nonparametric identification [^1] In this paper, “nonparametric identification” refers to identification under conditional independence constraints and additional functional restrictions, provided these are sufficiently weak in the sense that under relative support size/dimensionality assumptions on the variables, distribution violating them can be approximated arbitrarily well by distributions that satisfy them (Cui et al., 2024; Canay et al., 2013). of causal effects: (1) the bridge equation approach first introduced by Miao et al. (2018); and (2) the array decomposition approach, introduced to the causal literature by Kuroki and Pearl (2014) and Allman et al. (2015), building on ideas for full law identification in latent variable models which trace back to Kruskal (1977).
@@
The bridge equation approach was developed for settings in which proxies are available for an unmeasured confounder, allowing the counterfactual distribution to be recovered through re-expression of the standard confounding adjustment formula, without identification of the full law (Miao et al., 2018). The method leverages solutions to bridge equations—a generalization of solutions to linear systems—to recover the causal target. This method was later adapted to settings where proxies are available for an unmeasured mediator (Ghassami et al., 2025), recovering causal effects through re-expression of Judea Pearl’s front-door adjustment formula (Pearl, 1995).
@@
The array decomposition approach imposes invertibility conditions similar to those that ensure solutions to bridge equations, but rather than solving directly for a causal target, it recovers the full joint law of latent and observed variables, from which causal effects may be derived. Originating from the work of Kruskal (1977), this framework facilitates causal effect identification in models with key hidden variables beyond confounders, such as in settings with unmeasured exposures (Zhou and Tchetgen Tchetgen, 2024).
Both approaches identify causal effects in the presence of unmeasured confounding under distinct, non-nested conditional independence structures (Deaner, 2023). This naturally raises the question of how the frameworks compare under shared conditional independences, in terms of additional assumptions imposed.
This paper aims to synthesize and explicate the two frameworks, placing them on common footing under shared conditional independences. Since the assumptions underlying each framework are technically subtle and often difficult to interpret, clarifying their precise relationship is essential for both methodological development and applied use. Even among leading contributors to the field, there remains active discussion about how the two approaches differ, underscoring the need for careful exposition.
The assumptions underlying much of the existing literature are most transparent in fully discrete settings. In this regime, conditional distributions can be represented as matrices, and identification reduces to invertibility or rank conditions on these finite-dimensional objects. This algebraic perspective provides both technical clarity and intuition. Both frameworks, however, extend to models with continuous variables.
Accordingly, we frequently reference the discrete formulation throughout the paper to build intuition, and then discuss the continuous analogues in detail. Section 1 provides a brief overview of relevant background. In Section 2, we use a matrix representation to elucidate the assumptions underlying the bridge equation approach and to motivate their continuous analogues. Section 3 begins with the discrete formulation of the array decomposition approach, and then transitions to full law identification in continuous settings as developed by Hu and Schennach (2008) and extended to causal targets by Deaner (2023). Section 4 compares the two approaches under shared conditional independences in the presence of hidden confounding. Section 5 briefly discusses estimation strategies, and Section 6 concludes with directions for future work.
@@ 1.1 Causal Graphical Models
Bayesian networks, represented as directed acyclic graphs (DAGs), offer a structured framework for statistical modeling, where nodes denote random variables and edges encode conditional independence assumptions (Pearl, 1988). DAGs may be interpreted causally (Pearl, 2009; Spirtes et al., 2001), encoding counterfactual relationships which enable formal reasoning about causal effects and their identifiability from observed data (Richardson and Robins, 2013). For example, Figure 1 depicts a simple confounding structure in which a common cause $C$ influences both treatment $A$ and outcome $Y$. This structure entails the counterfactual independence statement $Y(a)\perp\!\!\!\perp A\mid C$, known as conditional ignorability, where $Y(a)$ denotes a potential outcome random variable, interpreted to mean “$Y$ had $A$, possibly contrary to fact, been set to value aa(Pearl, 2009).
The methods we discuss rely on conditional independences and counterfactual assumptions reflected in causal graphical models referenced throughout the paper. For an introduction to causal DAGs sufficient to understand this paper, see Elwert (2013).
@@ 1.2 Adjustment Formula
Consider the causal graphical model in Figure 1, where variables $C$ creating treatment-outcome ($A$ and $Y$, respectively) confounding are observed. The counterfactual distribution $f_{Y(a)}(y)$ is identified by the standard confounding adjustment formula,
$f_{Y(a)}(y)=\int f_{Y\mid A,C}(y\mid a,c)\,f_{C}(c)\,dc,$ (1)
where we use $f$ to denote both full and observed data distributions throughout, provided the following conditions hold for each $a\in\mathcal{A}$:
@@
In general, Assumption 1 holds if, by the rules of d-separation, conditioning on $C$ blocks all backdoor paths (which start with an arrowhead into $A$) from $A$ to $Y$ (Pearl, 2009; VanderWeele and Shpitser, 2013; Richardson and Robins, 2013).
@@
We provide an analogous review of the front-door adjustment formula from Pearl (1995) in Appendix B.
@@ 2 Proximal Causal Learning via Bridge Equations
We now present the assumptions of the proximal causal learning approach via bridge equations, introduced by Miao et al. (2018) as a framework for addressing hidden confounding using confounder proxies, recovering the counterfactual distribution via the standard adjustment formula. Ghassami et al. (2025) develop analogous assumptions leveraging proxies of a hidden mediator, enabling identification via the frontdoor adjustment formula in Equation 10, Appendix B.
If variables $C$ creating treatment-outcome confounding in Figure 1 are unobserved, the counterfactual distribution $f_{Y(a)}(y)$, and average causal effects, are not identifiable without further assumptions.
Consider the causal graphical model in Figure 2. In this figure, dashed edges ( ) denote optional causal relationships; their presence or absence does not change the results discussed in this paper. The variables $Z$ and $W$ are descendants of, and reasonably associated with, the unmeasured confounder $U$; we refer to them as proxies of $U$. Specifically, $Z$ is a treatment-inducing proxy, influencing $A$, while $W$ is an outcome-inducing proxy, influencing $Y$.
We emphasize that the term proxy is used informally and lacks a universally consistent definition. Alternatively, under the bridge equation approach, $Z$ may be a cause of $U$, $W$ may be a cause of $U$, or $A$ may be a cause of $Z$. Admissible configurations of directed acyclic graphs for this approach are described in Tchetgen Tchetgen et al. (2024).
Had $U$ been observed, the adjustment formula would yield the identifying expression from Equation 1, controlling for $U$ in place of $C$. Miao et al. (2018) replace the right hand side of Equation 1 with a functional of the observed law $f_{A,Y,W,Z}(a,y,w,z)$, leveraging Assumptions 4 and 7.
@@
Assumptions 4 and 5 are conditional independences reflected in the graphical model depicted in Figure 2. Miao et al. (2018) impose two additional conditions (Assumptions 6 and 7). We will elaborate upon these additional two assumptions and their meanings in Sections 2.1 and 2.2 below. A detailed proof of identification is given in their paper.
@@
It is important to emphasize that while this approach identifies the interventional distribution $f_{Y(a)}(y)$, it does not, in general, identify functionals that depend on UU—for instance, the conditional distribution $f_{Y\mid U,A}(y\mid u,a)$ or the marginal distribution $f_{U}(u)$ of the unobserved variable. This stands in contrast to the array decomposition approach discussed in Section 3.
2.1 Completeness Conditions
Assumption 6 is a generalized notion of linear independence in finite-dimensional vector spaces. To illustrate this idea, suppose for the time being that the model depicted in Figure 2 is fully discrete. Then, Assumption 6 means that the conditional distributions in $\{f_{Z\mid U,A}(z\mid u,a):u\in\mathcal{U}\}$ are a finite set of linearly independent vectors for each $a\in\mathcal{A}$. By Bayes’ rule, this is equivalent to linear independence of distributional vectors in $\{f_{U\mid Z,A}(u\mid z,a):u\in\mathcal{U}\}$ for each $a\in\mathcal{A}$.
To make the connection with linear algebra explicit, define $\mathbf{P}_{U\mid Z,a}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{Z}|}$ as in Appendix A with entries $\big(\mathbf{P}_{U\mid Z,a}\big)_{ij}\;:=\;\Pr(U=u_{i}\mid Z=z_{j},A=a).$ Then we have
@@ 2.1 Completeness Conditions
Hence, Assumption 6 holds if $\mathbf{P}_{U\mid Z,a}$ is right invertible. As noted in Cui et al. (2024), $Z$ must have at least as many categories as $U$ in order for Assumption 6 to hold.
Assumption 6 extends the notion of linear independence from finite-dimensional vector spaces to infinite-dimensional settings, where it is known as a completeness condition. Completeness conditions are notoriously abstruse, but many commonly used statistical families are known to satisfy Assumption 6. For example, the assumption holds when, for each $a\in\mathcal{A}$, the conditional distributions $\{f_{U\mid Z,A}(u\mid z,a):z\in\mathcal{Z}\}$ form an exponential family indexed by $z$, provided that $z$ has support on an open interval of the real line (Brown, 1986; Newey and Powell, 2003). At first, this formulation may appear more natural when $Z$ is viewed as causing $U$. However, if the family $\{f_{U\mid Z,A}(u\mid z,a):z\in\mathcal{Z}\}$ is specified as a regular exponential family indexed by $z$, then the corresponding conditional distributions $\{f_{Z\mid U,A}(z\mid u,a):u\in\mathcal{U}\}$ belong to the associated conjugate exponential family under standard conjugacy relationships.
Hu and Shiu (2022) provide additional characterizations of nonparametric families of complete distributions, including settings in which $U$ is generated as an arbitrary function of $Z$ with independent additive noise. In general, Assumption 6 is not verifiable (Canay et al., 2013). However, as long as $Z$ has larger dimension than $U$, any distribution violating Assumption 6 can be approximated arbitrarily well by distributions that satisfy them (Cui et al., 2024; Canay et al., 2013).
2.2 Solution to a Bridge Equation
@@
The integral equation in Assumption 7, referred to as a bridge equation, enables recovery of the counterfactual distribution $f_{Y(a)}(y)$ alongside prior assumptions. Since the existence of a solution is usually assumed point blank, leaving researchers no way to reason about when this assumption is reasonable, we unpack sufficient conditions for and implications of this assumption.
@@
In a fully discrete model, encode conditional distributions in matrices $\mathbf{P}_{Y\mid Z,a},\mathbf{P}_{W\mid Z,a},\mathbf{P}_{Y\mid U,a}$, and $\mathbf{P}_{W\mid U}$ (for details, see Appendix A). In this case, Assumption 7 is equivalent to the existence of a matrix $\mathbf{H}$ such that
$\mathbf{P}_{Y\mid Z,a}\;=\;\mathbf{H}\,\mathbf{P}_{W\mid Z,a}.$ (4)
A new completeness condition, Assumption 7a, is sufficient for Assumption 7 in the discrete setting.
Assumption 7a.
For each $a\in\mathcal{A}$ and every square-integrable function $g$,
$\mathbb{E}[g(Z)\mid W,A=a]=0\ \text{a.s.}\quad\Longleftrightarrow\quad g(Z)=0\ \text{a.s.}$
In the fully discrete case, this condition implies that $\mathbf{P}_{W\mid Z,a}$ admits a left inverse $\mathbf{P}_{W\mid Z,a}^{\dagger}$, so that
$\mathbf{P}_{Y\mid Z,a}=\mathbf{P}_{Y\mid Z,a}\,\mathbf{P}_{W\mid Z,a}^{\dagger}\,\mathbf{P}_{W\mid Z,a},$ (5)
which yields $\mathbf{H}=\mathbf{P}_{Y\mid Z,a}\,\mathbf{P}_{W\mid Z,a}^{\dagger}$ in Equation 4. This left invertibility condition requires $W$ to have at least as many categories as $Z$.
Via the independences in Assumption 4, we can decompose Equation 5 as such:
$\mathbf{P}_{Y\mid U,a}\mathbf{P}_{U\mid Z,a}=\mathbf{P}_{Y\mid Z,a}\,\mathbf{P}_{W\mid Z,a}^{\dagger}\,\mathbf{P}_{W\mid U}\mathbf{P}_{U\mid Z,a}$ (6)
Under Assumption 6, the matrix $\mathbf{P}_{U\mid Z,a}$ is right invertible. Consequently,
@@
Taking the inner product of both sides with the marginal vector $f_{U}$, the left-hand side yields entries $f_{Y(a)}(y)=\sum_{u}f_{Y\mid U,A}(y\mid u,a)\,f_{U}(u)$ and the right-hand side no longer relies on $u$ (since $\sum_{u}f_{W\mid U}(w\mid u)f_{U}(u)=f_{W}(w)$).
In continuous models, Assumptions 7b7d in addition to Assumption 7a are sufficient to ensure that Assumption 7 holds (Miao et al., 2018).